OpenStudy (anonymous):

How do you solve the following equation by using the graphing method? 3x + 6y = 18 2x + 5y = 20

OpenStudy (anonymous):

i would change them into y=mx+b graph the lines

hero (hero):

y = 3 - x/2 y = 4 - 2x/5

OpenStudy (anonymous):

yup

OpenStudy (curry):

yes thats right

OpenStudy (katrinakaif):

First Convert both of these equations to y = mx + b format: 3x + 6y = 18 --> 6y = -3x + 18 y = -1/2x + 3 And...: 2x + 5y = 20 --> 5y = 2x + 20 y = 2/5x + 4 Now graph: 1) Y = -1/2x + 3 - Point on yintercept of 3..and then Go down 1 and right 2 (hence slope is -1/2) Keep going down one and right 2 until you create a long enough line.. Do the same with the second equation, and then when done graphing, they should intersect. The point or coordinates that they intersect on is your y and x value.

OpenStudy (anonymous):

oh ok thank u :)

hero (hero):

Katrina, your second equation is wrong

OpenStudy (katrinakaif):

Anytime =)

OpenStudy (amistre64):

you dont need to convert them into anything; the equation of a line is the equation of a line regardless of whatever form its in ...

hero (hero):

Why does Katrina get three medals for getting it wrong?

hero (hero):

Tough crowd

OpenStudy (amistre64):

oh, it wasnt wrong :) just alot of extra work in my book

OpenStudy (amistre64):

well, apart from the wrongness lol

OpenStudy (amistre64):

if im gonna convert ill do this :) 3x + 6y = 18 x y --- + --- = 1 ; (0,3) (6,0) 6 3 2x + 5y = 20 x y --- + --- = 1 ; (0,4) (10,0) 10 4

OpenStudy (anonymous):

3x + 6y = 18 x intercept=(6,0) y intercept=(0,3 ) connect two points

hero (hero):

She asked about "graphing" it

OpenStudy (amistre64):

in order to graph, you gotta get at least 2 points per linear equation

hero (hero):

That's the reason why I always put it in y = mx + b form when I'm solving. I always graph it to check my answer

OpenStudy (katrinakaif):

She said how do you SOLVE BY graphing it..

OpenStudy (anonymous):



OpenStudy (anonymous):

do the same with other equation

OpenStudy (amistre64):



hero (hero):

Yeah, but she wanted to graph it on a calculator, using the "input" function....you know, using y = mx + b form

hero (hero):

You would have had to be here earlier to understand